OpenStudy (anonymous):

verify the identity: cos theta/1-sin theta = 1+sin theta/cos theta

OpenStudy (anonymous):

well solve the one on the left replace cos theta with (1-sin theta) and 1- sin theta to cos theta.

OpenStudy (anonymous):

oh no!

OpenStudy (anonymous):

that gives you the same identity to the right

OpenStudy (anonymous):

oh it should be 1+ sin theta

OpenStudy (anonymous):

hahahahahahahaha

OpenStudy (anonymous):

yeah except that \(\cos(\theta)\neq 1+\sin(\theta)\)

OpenStudy (anonymous):

\[\cos^2(\theta)=1-\sin^2(\theta)\\ \cos^2(\theta)=(1+\sin(\theta))(1-\sin(\theta))\] go from there

OpenStudy (anonymous):

thank you very much :)

OpenStudy (anonymous):

yw